two mechanics worked ona car the first mechanic charged 115 per hour and the second mechanic charged 65 per hour the mechanics worked for a combined total of 15 hours and together they charged a total of 1475 how long did each mechanic work
Answer by stanbon(75887)

two mechanics worked on a car. the first mechanic charged 115 per hour and the second mechanic charged 65 per hour the mechanics worked for a combined total of 15 hours and together they charged a total of 1475 how long did each mechanic work
Time Equation: f + s = 15 hrs
Cost Equation: 115f+65s = 1475 dollars
-------------------------------------------
Modify for elimination:
115f + 115s = 115*15
115f + 65s = 1475
-------
Subtract and solve for "s":
50s = 250
second mechanic = 5 hrs
------
Substitute for "s" and solve for "f"::
first mechanic = 15-5 = 10 hrs
------------
Cheers,
Stan H.
------------